OpenStudy (anonymous):

find amount of money in bank account if initial deposit is 4000 with annual rate of 2.5% for 5 years

OpenStudy (anonymous):

\[\huge A=P(1+\frac{r}{n})^{nt}\] P is principal, r is rate, n is number of times interest is compounded, t is time in years

OpenStudy (anonymous):

what is the n for

OpenStudy (anonymous):

is the interest compounded annually?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so we will have n=1

OpenStudy (anonymous):

\[\huge A=P(1+r)^t\]

OpenStudy (anonymous):

A=4000(1+2.5)^5 like that???

OpenStudy (anonymous):

\[\huge A=4000(1+.025)^5 \approx $4525.63\]

OpenStudy (anonymous):

ooh okay thanks